Nonlinear Retrieval of Wind Field via Optimal Technique

Hao Cheng, Guoqing Liu ^{†} and Wenzhong Ge ^{‡} †School of Sciences Nanjing University of Technology Nanjing, China 210009

chenghao95@yahoo.com.cn,guoqing@dns.njuct.edu.cn

‡

Department of Atmospheric Science Nanjing University Nanjing, China 210097 wzge@pub.jlonline.com

Abstract

In this paper, we develop a novel approach to nonlinearly retrieving wind field from radial wind field data, which are obtained by Doppler radar. In general, it is impossible to uniquely recover the 3-D wind field, since the problem is underdetermined.In order to overcome the difficulty, according to the real-life of the meteorology, we provide the constraints related with the piecewise smooth assumption of the wind field function for the wind retrieval.The corresponding nonlinear optimization can be solved by a modified affine scaling methodology, which is at a lower computational cost due to the behavior of the affine scaling technique. The real experiments demonstrate that the algorithms performs well.